1. Read the following and answer questions that follow.

The following information was taken from Masego Leero’s books.

01/01/2015 Started business with P20, 000 cash

03/01/2015 Deposited P15, 000 cash into the bank account

05/01/2015 Bought furniture on credit P3, 000 from Supreme

07/01/2015 Bought goods P7, 000 paying by cheque.

09/01/2015 Sold goods P6, 000 by cash

11/01/2015 Sold goods on credit to Lesego R P1, 000, Lorato L P2, 000 and Kago M P3, 000

13/01/2015 Kago M returned goods P300, Lorato L returned goods P200 to the business

15/01/2015 Bought goods from Sefalana worth P5, 000 and paid P3, 000 deposit by cheque

17/01/2015 Returned P500 goods to Sefalana

19/01/2015 Paid a cab for transporting goods in the business P300 by cash.

21/01/2015 Lesego R paid us P500 cash, Lorato L and Kago M paid us the amount owed by cheque

23/01/2015 Bought goods from the following on credit: Choppies P3, 000, Superspar P5,000

25/01/2015 Sold goods by cheque P10, 000

26/01/ 2015 Bought a second hand car paying by cheque P8, 000

29/01/2015 Paid electricity P500, telephone P500, rent P2, 000 by cheque

30/01/2015 Bought stationary by cash P300

31/01/2015 Paid Choppies P2, 000, Superspar P3, 000 by cheque

31/01/2015 Sold goods by cheque P3, 000

e. Using the new trial balance and additional notes given below:

Notes:

1. Closing inventory on 31 January 2015 was P2, 000

2. Assets are to be depreciated at 25% at cost on a reducing balance method

3. Interest payable P100 was owing

4. Electricity P100 was paid in advance

5. Provision should be made for bad debts at 10% of debtors

i. Draw up the income statement (11 marks)

ii. Statement of Financial Position (8 marks)

f. Use the following ratios to analyse and interpret the results of the financial statements. (10 marks)

i. Gross profit margin (2 marks)

ii. Gross profit mark up (2 marks)

iii. Stock turn over (2marks)

vi. Current ratio (2 marks)

vii. Acid test ratio (2 marks)
Answer #1

Income statement

To Purchase (goods) 7000 By sales (goods) 19000
To purchase By sales (goods) Credit 6000
Seflana 5000 Less Returned 500 5500
less returned 500 Closing stock 2000
Choppies 3000
Superspar 5000 12500
To taxi charges 300
To electricity(500-100 prepaid) 400
To Telephone 500
To rent 2000
To stationery expense 300
To Depreciation 2750
To interest Expense 100
To Bad debts 50
To Net Profit 600
Total 26500 Total 26500

Financial Statement (Balance sheet)

Capital 20000 Cash 10900
Add Profit 600 Bank a/c 6500
Accounts Payable( less returned) Furniture (3000 - 750 Depreciation @ 25% RBM) 2250
Supreme 3000 Accounts Receivable(less Returned goods)
Seflana 2000-500 Lesego R 500 500
   Superspar 2000 7500   
Choppies 1000 Second hand Car (8000- 2000 Depreciation @ 25% RBM)    6000
Accrued Interest 100 Closing Stock 2000
Provision for Bad debt 50 Prepaid Electricity 100
28250 28250

Gross Profit Margin Ratio= Gross Profit/Sales *100

Gross Profit=Sales + Closing stock - Purchases

= 27500-19500

=7000

=7000/25500*100

=27.45%

Gross Profit markup= 27%

Stock turnover ratio = Closing Stock/sales*100

=2000/25500*100

=7.84%

Current ratio=Current assets/current liabilities

=All current assets except car and furniture=20000

Current liabilities= All current liabilities except Capital and profit=7650

=20000/7650

=2.61

Acid Test Ratio=Liquid assets/Liquid liabilities

= Cash and bank/Current liabilities

=17400/7650

=2.274

Looking at the ratios we can interpret that the company is enjoying a strong financial position as its assets are enough that is twice than its liabilities and liquidity is also sound and it has enough funds to repay its liabilities.